OpenStudy (anonymous):

Check all polynomials that are perfect square trinomials: a. 49x^2 − 8x + 16 b. 4a^2 − 20a + 25 c. 25b^2 − 5b + 10 d. 16x^2 − 8x + 1

OpenStudy (anonymous):

b , d

OpenStudy (anonymous):

4a^2 − 20a + 25 = (2x - 5)^2 16x^2 − 8x + 1 = (4x - 1)^2

OpenStudy (anonymous):

The patterns you are looking for are\[a^2+2ab+b^2=(a+b)^2\]or\[a^2-2ab+b^2=(a-b)^2\]